# MADAR: Efficient Continual Learning for Malware Analysis with Diversity-Aware Replay

### v1

@@ Abstract
Millions of new pieces of malicious software (i.e., malware) are introduced each year. This poses significant challenges for antivirus vendors, who use machine learning to detect and analyze malware, and must keep up with changes in the distribution while retaining knowledge of older variants. Continual learning (CL) holds the potential to address this challenge by reducing the storage and computational costs of regularly retraining over all the collected data. Prior work, however, shows that CL techniques, which are designed primarily for computer vision tasks, fare poorly when applied to malware classification. To address these issues, we begin with an exploratory analysis of a typical malware dataset, which reveals that malware families are diverse and difficult to characterize, requiring a wide variety of samples to learn a robust representation. Based on these findings, we propose $\underline{M}$alware $\underline{A}$nalysis with $\underline{D}$iversity-$\underline{A}$ware $\underline{R}$eplay (MADAR), a CL framework that accounts for the unique properties and challenges of the malware data distribution. Through extensive evaluation on large-scale Windows and Android malware datasets, we show that MADAR significantly outperforms prior work. This highlights the importance of understanding domain characteristics when designing CL techniques and demonstrates a path forward for the malware classification domain.
Index Terms:
Malware Analysis; Windows Malware; Android Malware; Catastrophic Forgetting; Continual Learning;
@@ I Introduction
Advances in machine learning have significantly enhanced the detection and classification of malicious software, achieving notable success across various domains such as Windows executables [1, 2, 3], PDFs [4], and Android applications [5, 6]. Traditional models, trained on static datasets, are typically expected to perform well on new data under the assumption of a constant data distribution. However, in reality, both malicious (malware) and benign (goodware) software evolve continuously, necessitating regular model updates to adapt to changes in data distribution and maintain effectiveness. For example, the AV-TEST Institute reports approximately 450,000 new malware samples daily [7], while VirusTotal processes over one million unique submissions each day [8]. This scale creates immense challenges in training and even storing all the samples being identified.
@@
Training a malware classification model solely on new data can lead to catastrophic forgetting [9, 10], where previously learned information is lost, resulting in misclassification of benign software or allowing attackers to evade detection with older malware strains, known as Retrograde Malware Attacks (RMA) (see Section II). As such, anti-virus vendors must deal with difficult trade-offs: (i) removing older samples from the training set, risking exposure to revived older malware; (ii) ignoring newer samples, risking failure to detect emerging trends; (iii) reducing the frequency of retraining, compromising accuracy during intervals between updates; or (iv) incurring significant effort and cost to frequently retrain on the entire dataset. These challenges highlight the need for agile and adaptive malware classification techniques capable of learning incrementally and responding to the dynamic malware landscape.
@@
Continual learning (CL) provides a promising solution to this problem by enabling models to adapt to new data without requiring the retention of large datasets or frequent retraining [11, 12, 13, 14, 15, 16]. By addressing catastrophic forgetting, CL techniques ensure that models remain effective and efficient in the face of evolving malware distributions. While designs for CL have been extensively studied in the context of computer vision [17, 18, 14], there are very few such studies in the malware classification domain [19, 11]. Rahman et al. [19] observed that CL techniques originally developed for computer vision problems fail to deliver acceptable performance in malware classification, due in part to the strong semantics of malware features and the high level of diversity found in the malware ecosystem.
@@
In this study, we first delve into the complexities of malware data distributions using the EMBER dataset [20] of Windows malware and goodware. Our analysis highlights the diversity in malware, both between and even within families, or groups of related malware. Leveraging this insight, we devise MADAR – $\underline{M}$alware $\underline{A}$nalysis with $\underline{D}$iversity-$\underline{A}$ware $\underline{R}$eplay, a replay-based continual learning strategy that accounts for diversity and achieves improved malware classification performance. In particular, MADAR replays a mix of representative samples and novel samples (i.e., outliers) to enhance the model’s ability to retain knowledge and identify new malware variants despite memory constraints. Our techniques employ Isolation Forests (IF) [21] to identify critical novel samples, either directly through the raw feature vectors (MADAR), or through the use of model activations for a more compact representation (MADARθ). For both of these approaches, we also consider two mechanisms to control how the number of replay samples is chosen, which refer to as Ratio and Uniform strategies.
@@
We evaluate these techniques with comprehensive experiments on two large-scale datasets across three CL scenarios representative of real-world malware analysis tasks, covering domain incremental learning (Domain-IL), class incremental learning (Class-IL), and task incremental learning (Task-IL). These datasets include the well-known EMBER dataset [20], containing one million examples of Windows executables, and a new benchmark dataset of Android malware from the AndroZoo repository [22] specifically created to explore CL scenarios. Our experimental results on these datasets confirm that MADAR is indeed effective and much better than prior state-of-the-art CL methods in the face of realistic malware data distribution shifts.
@@
  • We provide an exploratory analysis of the diversity of malware distributions and show how it creates unique challenges for continuous learning.
@@ II Threat Model
The Retrograde Malware Attack (RMA) describes an attack scenario where adversaries can exploit catastrophic forgetting in machine learning-based malware detection systems. RMA occurs when a system, updated incrementally with only new samples, loses the ability to recognize older malware, allowing attackers to reintroduce legacy or slightly modified variants that evade detection.
@@ III-A Replay in Continual Learning
Exact Replay.  Selecting and utilizing replay samples in exact replay involves determining a memory budget, denoted as $\mathcal{M}$. Finding the optimal way to choose $\mathcal{M}$ remains an open research question [16, 23]. Exact-replay techniques are designed to choose replay samples from previously learned data to be combined with new samples for retraining. The goal of these techniques is to maximize the performance with minimal replay samples [24, 25, 15, 26].
@@
Compressed Replay.  Compressed replay, also known as latent replay, focuses on storing and replaying activations from an intermediate layer of a previous neural network to prevent forgetting. In classification tasks, these representations from hidden layers are replayed into selected layers based on the architecture and task at hand [28, 29, 30].
@@
Despite extensive work in CL, very few studies have ventured into applying CL in the realm of malware. To the best of our knowledge, Rahman et al.[19] were the first to explore CL for malware classification. They concluded that existing CL methods fall short in tackling forgetting in malware classification systems due to differences in the underlying nature of the data distribution shifts that occur in practice versus those explored in the computer vision domain. Malware representations leverage tabular features with strong semantic constraints that limit the space of feasible samples, and within that space, samples exhibit a high level of diversity. Replay-based techniques were found to perform better compared to other approaches in this setting. More recently, Chen et al. proposed to combine contrastive learning with active learning to continuously train Android malware classifiers [3]. They focus on the detection of concept drift, rather than overcoming forgetting.In addition, some studies have used online learning for malware classification, which deals with adding new samples as they are observed but does not directly address catastrophic forgetting [31]. Another CL domain in cybersecurity is network intrusion detection (NID), looking for malicious activity based on network packets. Amalapuram et al. explored a replay-based CL technique that incorporates class-balancing reservoir sampling and perturbation assistance for parameter approximation NID [32]. Another recent work explored semi-supervised CL for NID in a class incremental setting [33]. We note that NID is a different domain with different data characteristics than malware. Further, these works do not focus on reducing CF.
@@ IV-A CL Scenarios for Malware Classification
Task-IL
In malware analysis, leveraging insights from additional methods can prove beneficial. This may involve identifying the broader malware category (e.g., adware, ransomware, etc.), malware behaviors [37], or the infection vector (e.g., phishing, downloader, etc.). Task-IL encapsulates this concept of constrained tasks, where the introduction of a new task may symbolize a new category or behavior set. This event occurs less frequently than adding a new family, as seen in Class-IL, yet it poses a genuine issue in malware classification. Unlike Class-IL, the task identity is provided to the model at test time, significantly simplifying the problem. In malware, this could mean learning the task identity from a separate model, manual analysis, or field reports of the malware’s behavior. However, as our datasets do not possess naturally defined tasks, we partition our dataset into tasks comprising an equal number of independent and non-overlapping classes to act as a proxy to new behaviors, following common practice in the CL literature [34, 14]. In other words, a given task would be to perform family classification among one subset of families, and the subset that each sample belongs to is known to the classifier. The model is expected to be able to handle multiple tasks at once, and new tasks are being added during each experiment.
Figure 3: CL scenarios in a typical malware analysis pipeline.
@@ IV-B Dataset
In this study, we use large-scale Windows and Android malware datasets: EMBER [20], a Windows malware dataset from prior work, recognized as a standard benchmark for malware classification, and two new Android malware datasets derived from AndroZoo [22], specifically assembled for this research.
@@ IV-B1 Windows PE Files
For our experiments, we leverage the EMBER 2018 dataset, containing features from one million Windows Portable Executable (PE) files, predominantly scanned in 2018111https://github.com/elastic/ember. The dataset comprises 400K goodware and 350K malware, with the rest labeled as unknown. EMBER provides a diverse array of 2,381 hand-crafted features, covering general file information, header data, import/export functions, and section details. Notably, these features capture strong semantic concepts that have a limited space of feasible settings, outside of which the executable does not actually run.
@@ IV-B2 Android APK Files
Additionally, we collected two datasets from AndroZoo [22] (AZ) for our experiments: AZ-Domain for Domain-IL and AZ-Class for Class-IL and Task-IL. These datasets contain Android APK files, and both use a 9:1 ratio of goodware to malware to reflect real-world class imbalance. Following the practice of prior work [31], the malware samples are selected with a VirusTotal detection count of $>=4$. The AZ-Domain dataset includes 80,690 malware and 677,756 goodware samples from 2008 to 2016. We divided the AZ-Domain dataset into non-overlapping yearly training and testing sets. The AZ-Class dataset consists of 285,582 samples from 100 Android malware families, each with at least 200 samples.
We extracted Drebin features [5] from the apps for both datasets. These features cover various aspects of app behavior, including hardware access, permissions, app component names, filtered intents, restricted API calls, used permissions, suspicious API calls, and network addresses. Again, we note that these capture strong semantic concepts from the operation of the application. The training sets of AZ-Domain and AZ-Class have 3,858,791 and 1,067,550 features, respectively. We processed the test datasets to match the training feature sets and reduced dimensionality by filtering features with low variance ($<0.001$) using scikit-learn’s VarianceThreshold. This resulted in final feature dimensions of 1,789 for AZ-Domain and 2,439 for AZ-Class, respectively.
@@ IV-D Baselines and Metric
We adopt two baselines for comparison: None and Joint. None sequentially trains the model on each new task without any CL techniques, serving as an informal minimum baseline. By contrast, Joint uses all new and prior data for training at each step, acting as an informal maximum baseline. Despite its resource demands, Joint ensures strong performance throughout the dataset. We also introduce an additional baseline – Global Reservoir Sampling – which provides an unbiased sampling of the underlying class distributions and serves as a strong point of comparison for our diversity-aware approach.
@@ IV-E Training and Evaluation Protocol
In our evaluations, we use a non-overlapping hold-out set corresponding to each task. For example, the AZ-Domain dataset contains 8 years of training samples from 2008 to 2016, resulting in 9 hold-out sets, one for each year. A CL model is evaluated on all the hold-out sets up to the current task; formally, the model is evaluated on tasks $t_{i}$ to $t_{T}$, for $1\leq i\leq T$, after it been trained on the current task $t_{T}$.
In this work, each set of experiments is performed around 10-15 times with different random parameter initializations. We use PyTorch on a CentOS-7 machine with an Intel Xeon processor, 40 CPU cores, 128GB RAM, and four GeForce RTX 2080Ti GPUs, each with 12GB memory.
@@ V Exploratory Analysis of EMBER
$913091309130$JanFebMarAprMayJuneJulyAugSepOctNovDecMonthsFrequencyNew FamiliesLearned Families
@@
In addition, many malware families display complex distributional patterns in feature space, making for additional diversity within classes. Figure 6, for example, shows a t-SNE projection of the EMBER features for all malware samples from January 2018. Each class (represented by color) is not clustered into a single well-defined region. Rather, the larger classes are split up into subsets that spread out in feature space. To accurately represent the malware distribution, it is thus important to select samples not only from each class, but also from multiple areas within each class.
Despite some degree of overlap among different families, it is possible to discern distinct clusters indicative of the diversity across and even within malware families. It is vital to capture samples from each of these smaller clusters, including those belonging to minority families, to accurately represent the malware landscape for a specific task month. To further complicate the situation, it is often not possible to provide definitive family labels for a sample due to the inherent subjectivity involved in malware analysis, which results in a large, diverse set of additional unlabeled samples which must be considered [36]. These attributes of the malware landscape make it difficult to characterize classes as contiguous regions of the feature space with relatively simple class boundaries. Rather, each class might only be understood as a collection of smaller pockets of the feature space that might be closer to pockets from other classes than the same class. This may explain why prior CL techniques designed for computer vision datasets are less effective when applied to the malware domain [19].
@@
In light of these analysis, we propose that selecting samples based on families and variations within those families could more effectively capture the diversity within the replay data distribution, potentially mitigating catastrophic forgetting.
VI Diversity-Aware Replay
Here, we introduce the our proposed continual learning (CL) framework for malware classification with two diversity-aware replay variants: MADAR and MADARθ. MADAR operates on the basis of the raw-feature space and MADARθ operates on the basis of the model’s weights-space of raw-features.

### v2

# MADAR: Efficient Continual Learning for Malware Analysis with Distribution-Aware Replay

@@ Abstract
Millions of new pieces of malicious software (i.e., malware) are introduced each year. This poses significant challenges for antivirus vendors, who use machine learning to detect and analyze malware, and must keep up with changes in the distribution while retaining knowledge of older variants. Continual learning (CL) holds the potential to address this challenge by relaxing the requirements of the incremental storage and computational costs of regularly retraining over all the collected data. Prior work, however, shows that CL techniques, which are designed primarily for computer vision tasks, fare poorly when applied to malware classification. To address these issues, we begin with an exploratory analysis of a typical malware dataset, which reveals that malware families are heterogeneous and difficult to characterize, requiring a wide variety of samples to learn a robust representation. Based on these findings, we propose $\underline{M}$alware $\underline{A}$nalysis with $\underline{D}$istribution-$\underline{A}$ware $\underline{R}$eplay (MADAR), a CL framework that accounts for the unique properties and challenges of the malware data distribution. Through extensive evaluation on large-scale Windows and Android malware datasets, we show that MADAR significantly outperforms prior work. This highlights the importance of understanding domain characteristics when designing CL techniques and demonstrates a path forward for the malware analysis domain.
@@ I Introduction
Advances in machine learning have significantly enhanced the detection and classification of malicious software, achieving notable success across various domains such as Windows executables [1, 2, 3], PDFs [4], and Android applications [5, 6, 7]. Traditional models, trained on static datasets, are typically expected to perform well on new data under the assumption of a constant data distribution. However, in reality, both malicious (malware) and benign (goodware) software evolve continuously, necessitating regular model updates to adapt to changes in data distribution and maintain effectiveness. For example, the AV-TEST Institute reports approximately 450,000 new malware samples daily [8], while VirusTotal processes over one million unique submissions each day [9]. This scale creates immense challenges in training and even storing all the samples.
@@
Training a malware classification model solely on new data can lead to catastrophic forgetting (CF) [10, 11], where previously learned information is forgotten, resulting in increased misclassification and allowing attackers to evade detection with older malware strains, known as Retrograde Malware Attacks (RMA) (see Section II). As such, anti-virus vendors must deal with difficult trade-offs: (i) removing older samples from the training set, risking exposure to revived older malware; (ii) ignoring newer samples, risking failure to detect emerging trends; (iii) reducing the frequency of retraining, compromising accuracy during intervals between updates; or (iv) incurring significant effort and cost to frequently retrain on the combined new and older samples. These challenges highlight the need for agile and adaptive malware classification techniques capable of learning incrementally and responding to the dynamic malware landscape.
@@
Continual learning (CL) provides a promising solution to this problem by enabling models to adapt to new data without requiring the retention of large datasets or frequent retraining [12, 13, 14, 15, 16, 17]. By addressing catastrophic forgetting, CL techniques ensure that models remain effective and efficient in the face of evolving malware distributions. While designs for CL have been extensively studied in the context of computer vision [18, 19, 15], there are very few such studies in the malware classification domain [20, 12]. Rahman et al. [20] observed that CL techniques originally developed for computer vision problems fail to deliver acceptable performance in malware classification, due in part to the strong semantics of malware features and the high level of heterogeneity found in the malware ecosystem.
@@
In this study, we first delve into the complexities of malware data distributions using the EMBER dataset [21]  [^1] The recently released EMBER 2024 dataset [22] is outside the scope of this study, as it was not available at the time the experiments were conducted.. Our analysis highlights the heterogeneity in malware, both between and even within families, or groups of related malware. Leveraging this insight, we devise MADAR – $\underline{M}$alware $\underline{A}$nalysis with $\underline{D}$istribution-$\underline{A}$ware $\underline{R}$eplay, a replay-based continual learning strategy that accounts for heterogeneity and achieves improved malware classification performance. In particular, MADAR replays a mix of representative samples and novel samples (i.e., outliers) to enhance the model’s ability to retain knowledge and identify new malware variants despite memory constraints. Our techniques employ Isolation Forests (IF) [23] to identify critical novel samples, either directly through the raw feature vectors (MADAR), or through the use of the hidden representations of the model for a more compact representation (MADARθ). For both of these approaches, we consider two mechanisms to control how the number of replay samples is chosen, which we refer to as Ratio and Uniform strategies.
@@
We evaluate these techniques with comprehensive experiments on two large-scale datasets across three CL scenarios representative of real-world malware analysis tasks, covering domain incremental learning (Domain-IL), class incremental learning (Class-IL), and task incremental learning (Task-IL). These datasets include the well-known EMBER dataset [21], containing one million examples of Windows executables, and a new benchmark dataset of Android malware from the AndroZoo repository [24] specifically created to explore CL scenarios. Our experimental results across these datasets demonstrate that MADAR is effective and outperforms prior state-of-the-art continual learning methods when confronted with realistic distribution shifts in malware data.
@@
  • We provide an exploratory analysis of the heterogeneity of malware distributions and show how it creates unique challenges for continuous learning.
@@ II Threat Model
The Retrograde Malware Attack (RMA) describes an attack scenario where adversaries can exploit catastrophic forgetting in machine learning-based malware detection systems. RMA occurs when a system, updated incrementally with only new samples, loses the ability to recognize older malware, allowing attackers to reintroduce legacy or slightly modified variants that can evade detection.
@@ III-A Replay in Continual Learning
Exact Replay.  Selecting and utilizing replay samples in exact replay involves determining a memory budget, denoted as $\mathcal{B}$. Finding the optimal way to choose $\mathcal{B}$ remains an open research question [17, 25]. Exact-replay techniques are designed to choose replay samples from previously learned data to be combined with new samples for retraining. The goal of these techniques is to maximize the performance with minimal replay samples [26, 27, 16, 28].
@@
Despite extensive work in CL, very few studies have ventured into applying CL in the realm of malware. To the best of our knowledge, Rahman et al.[20] were the first to explore CL for malware classification. They concluded that existing CL methods fall short in tackling forgetting in malware classification systems due to differences in the underlying nature of the data distribution shifts that occur in practice versus those explored in the computer vision domain. Malware representations leverage tabular features with strong semantic constraints that limit the space of feasible samples, and within that space, samples exhibit a high level of heterogeneity. Replay-based techniques are found to perform better compared to other approaches in this setting. Another recent work, MalCL [12], introduces a generative replay-based continual learning approach for malware classification using a generative adversarial network (GAN). It achieves state-of-the-art performance compared to prior generative replay methods. However, its evaluation is restricted to the Class-IL scenario.
Furthermore, Chen et al. proposed to combine contrastive learning with active learning to continuously train Android malware classifiers [3]. They focus on the detection of concept drift, rather than overcoming forgetting.In addition, some studies have used online learning for malware classification, which deals with adding new samples as they are observed but does not directly address catastrophic forgetting [30]. Another CL domain in cybersecurity is network intrusion detection (NID), looking for malicious activity based on network packets. Amalapuram et al. explored a replay-based CL technique that incorporates class-balancing reservoir sampling and perturbation assistance for parameter approximation NID [channappayya2024augmented]. Another recent work explored semi-supervised CL for NID in a class incremental setting [31]. We note that NID is a different domain with different data characteristics than malware. Further, these works do not focus on reducing CF.
@@ IV-A CL Scenarios for Malware Classification
Figure 3: CL scenarios in a typical malware analysis pipeline.
@@ IV-A2 Class-IL
IV-A3 Task-IL
In malware analysis, leveraging insights from additional methods can prove beneficial. This may involve identifying the broader malware category (e.g., adware, ransomware, etc.), malware behaviors [35], or the infection vector (e.g., phishing, downloader, etc.). Task-IL encapsulates this concept of constrained tasks, where the introduction of a new task may symbolize a new category or behavior. This event occurs less frequently than adding a new family, as seen in Class-IL, yet it poses a genuine issue. Unlike Class-IL, the task identity is provided to the model at test time, significantly simplifying the problem. In malware, this could mean learning the task identity from a separate model, manual analysis, or field reports of the malware’s behavior. However, as our datasets do not possess naturally defined tasks, we partition our dataset into tasks comprising an equal number of independent and non-overlapping classes to act as a proxy to new behaviors, following common practice in the CL literature [32, 15]. In other words, a given task would be to perform family classification among one subset of families, and the subset that each sample belongs to is known to the classifier. The model is expected to be able to handle multiple tasks at once, and new tasks are being added during each experiment.
@@ IV-B Dataset
In this study, we use large-scale Windows and Android malware datasets: EMBER [21], a Windows malware dataset from prior work, recognized as a benchmark for malware classification, and two new Android malware datasets derived from AndroZoo [24], specifically assembled for this research.
@@ IV-B1 Windows PE Files
For our experiments, we leverage the EMBER 2018 dataset, containing features from one million Windows Portable Executable (PE) files, predominantly scanned in 2018  [^2] https://github.com/elastic/ember. The dataset comprises 400K goodware and 350K malware, with the rest labeled as unknown. EMBER provides a diverse array of 2,381 hand-crafted features, covering general file information, header data, import/export functions, and section details. Notably, these features capture strong semantic concepts that have a limited space of feasible settings, outside of which the executable does not actually run.
@@ IV-B2 Android APK Files
Additionally, we collected two datasets from AndroZoo [24] (AZ) for our experiments: AZ-Domain for Domain-IL and AZ-Class for Class-IL and Task-IL. These datasets contain Android APK files, and both use a 9:1 ratio of goodware to malware to reflect real-world class imbalance. Following the practice of prior work [30], the malware samples are selected with a VirusTotal detection count of >=4>=4. The AZ-Domain dataset includes 80,690 malware and 677,756 goodware samples from 2008 to 2016. We divided the AZ-Domain dataset into non-overlapping yearly training and testing sets. The AZ-Class dataset consists of 285,582 samples from 100 Android malware families, each with at least 200 samples.
We extracted Drebin features [5] from the apps for both datasets. These features cover various aspects of app behavior, including hardware access, permissions, app component names, filtered intents, restricted API calls, used permissions, suspicious API calls, and network addresses. Again, we note that these capture strong semantic concepts from the operation of the application. The training sets of AZ-Domain and AZ-Class have 3,858,791 and 1,067,550 features, respectively. We processed the test datasets to match the training feature sets and reduced dimensionality by filtering features with low variance (<0.001<0.001) using scikit-learn’s VarianceThreshold. This resulted in final feature dimensions of 1,789 for AZ-Domain and 2,439 for AZ-Class, respectively.
@@ IV-D Baselines and Metric
We adopt two baselines for comparison: None and Joint. None sequentially trains the model on each new task without any CL techniques, serving as an informal minimum baseline. By contrast, Joint uses all new and prior data for training at each step, acting as an informal maximum baseline. Despite its resource demands, Joint ensures strong performance throughout the dataset. We also introduce an additional baseline – Global Reservoir Sampling – which provides an unbiased sampling of the underlying class distributions and serves as a strong point of comparison for our distribution-aware approach.
@@ IV-E Training and Evaluation Protocol
In our evaluations, we use a non-overlapping hold-out set corresponding to each task. For example, the AZ-Domain dataset contains 8 years of training samples from 2008 to 2016, resulting in 9 hold-out sets, one for each year. A CL model is evaluated on all the hold-out sets up to the current task; formally, the model is evaluated on tasks $t_{i}$ to $t_{T}$, for $1\leq i\leq T$, after it been trained on the current task $t_{T}$. In this work, each set of experiments is performed around 10-15 times with different random parameter initializations. We use PyTorch on a CentOS-7 machine with an Intel Xeon processor, 40 CPU cores, 128GB RAM, and four GeForce RTX 2080Ti GPUs, each with 12GB memory.
@@ V Exploratory Analysis of EMBER
913091309130913091309130913091309130913091309130JanFebMarAprMayJuneJulyAugSepOctNovDecMonthsFrequencyNew FamiliesLearned Families
@@
In addition, many malware families display complex distributional patterns in feature space, making for additional heterogeneity within classes. Figure 6, for example, shows a t-SNE projection of the EMBER features for all malware samples from January 2018. Each class (represented by color) is not clustered into a single well-defined region. Rather, the larger classes are split up into subsets that spread out in feature space. To accurately represent the malware distribution, it is thus important to select samples not only from each class, but also from multiple areas within each class.
Despite some degree of overlap among different families, it is possible to discern distinct clusters indicative of the heterogeneity across and even within malware families. It is vital to capture samples from each of these smaller clusters, including those belonging to minority families, to accurately represent the malware landscape for a specific task month. To further complicate the situation, it is often not possible to provide definitive family labels for a sample due to the inherent subjectivity involved in malware analysis, which results in a large, diverse set of additional unlabeled samples which must be considered [34]. These attributes of the malware landscape make it difficult to characterize classes as contiguous regions of the feature space with relatively simple class boundaries. Rather, each class might only be understood as a collection of smaller pockets of the feature space that might be closer to pockets from other classes than the same class. This may explain why prior CL techniques designed for computer vision datasets are less effective when applied to the malware domain [20].
@@
In light of these analysis, we propose that selecting samples based on families and variations within those families could more effectively capture the heterogeneity within the replay data distribution, potentially mitigating catastrophic forgetting.
VI Distribution-Aware Replay
Here, we introduce the our proposed continual learning (CL) framework for malware classification with two distribution-aware replay variants: MADAR and MADARθ. MADAR operates on the basis of the raw-feature space and MADARθ operates on the basis of the model’s weights-space of raw-features.